# Morpheus: Benchmarking Physical Reasoning of Video Generative Models with Real Physical Experiments

### v1

@@ Appendix A Dataset
Dataset Description: The dataset comprises recordings of five distinct real-world dynamic systems: a falling ball, a non-holonomic pendulum, a holonomic pendulum, and a projectile motion. For each system, we recorded multiple times the type of experiment trying to have homogenous videos, while after a few iterations, we varied the initial conditions or configuration parameters. Table A1 below summarizes the number of recordings and configurations for each experiment.
Experiment Videos Factors of Variation Configuration/ Initial Condition Description
Falling Ball 12 1 Height from which the ball was released.
@@
Non-Holonomic Pendulum 15 1 Initial angle from the vertical (zero-degree resting position).
@@
Table A1: Summary of the Experimental Dataset from real-world recorded videos.
Falling ball
For the falling ball experiment, we used a normal table tennis orange ball. A mechanic actuator  [^2] Motor Model: T825, Motor serial number: 00362129 was used to hold the ball at a certain height (initial position) and as a release mechanism to control the moment the ball was let at a free fall, before making contact with the surface below. Different height levels from the surface were used as initial positions, resulting in trajectories with different lengths (smaller or larger).
Figure A1: Representative frames of the falling ball experiment. A mechanical actuator releases the ball from different heights. The final frame of each sequence marks the exact moment the ball contacts the surface.
@@
Figure A2: Representative frames of the bouncing ball experiment. A mechanical actuator releases the ball from different heights. This experiment begins immediately after the ball’s initial contact with the surface and ends after the next touch.
@@
Figure A3: Representative frames of the projectile motion experiment with a 3D-printed launcher. In the dataset the launch angle, force, and ball color are varied.
@@
Figure A4: Representative frames of the holonomic pendulum experiment. The pendulum is released by a mechanical actuator and swings from varying starting positions, resulting in different initial angles.
Non-holonomic pendulum
The setup for the non-holonomic pendulum is almost identical to that of the holonomic one, with the difference being that instead of a non-deformable metal structure, a flexible metal string was employed. The metal string on its top end had a noose so that it can be hanged from certain points, while at the other end, we attached a standard table tennis ball as in the holonomic pendulum case. Using the same release mechanism we can control the initial angular position with respect to the resting (vertical) position, at the very beginning of the trajectory, while the object is still stationary.
Figure A5: Representative frames of the non-holonomic pendulum experiment. A flexible metal string suspends a table tennis ball, with initial angles controlled by a mechanical actuator functioning as release mechanism.
@@
Figure A6: Representative frames of the double pendulum experiment. A custom-built structure with a metal pole and joint supports the 3D-printed plastic bobs, allowing for free rotation with minimal friction. The initial angles of both pendulum links are precisely controlled using a mechanical actuator.
Appendix B Velocity and acceleration estimation
@@
$\begin{bmatrix}v\\ b\end{bmatrix}=\left(A^{T}A\right)^{-1}A^{T}x$ (A3)
@@
$A=\begin{bmatrix}t_{1}&1\\ t_{2}&1\\ \vdots&\vdots\\ t_{w}&1\end{bmatrix}$ (A4)
@@
$v_{\text{final}}=\alpha v_{\text{regression}}+(1-\alpha)v_{\text{central}}$ (A5)
Here $\alpha=0.7$, indicating greater confidence in the regression method. Finally, we apply Savitzky-Golay filtering for smoothing [37]. This step effectively removes high-frequency noise from velocity calculations.
$v_{\text{smoothed}}=\text{SG}(v_{\text{final}},\text{window},3)$ (A6)
@@
$v(t)=\text{SG}\left(\alpha v_{\text{regression}}(t)+(1-\alpha)v_{\text{central}}(t),w,3\right)$ (A7)
where w is the window size (odd number for symmetry); $\alpha=0.7$ is the weighting coefficient; SG represents Savitzky-Golay filter of order 3; Regression window range: $[t-w/2,t+w/2]$.
For the acceleration, we first calculate the acceleration using the central difference. For $21\leq i\leq N-$ 2:
$a_{i}=\frac{v_{i+1}-v_{i-1}}{t_{i+1}-t_{i-1}}$ (A8)
@@
Appendix C Evaluation metrics
C.1 Discard rate
We generate $N_{total}$ videos for each type of experiment. Among these videos, we discard those that do not meet our quality standards, following a three-stage filtering out. First, we discard videos where object are disappearing from the videos the number of such videos is $N_{disappear}$. Second, we analyze the number of objects in each video and discard videos that do not maintain a consistent object count in the not discarded yet videos. For this purpose we employ DEVA tracking [38] built on top of Grounded SAM [39] (with object names from the prompt as Grounding DINO [40] query) for consistent open-vocabulary prediction of 2D object masks. We denote the number of discarded videos in this step as $N_{duplicate}$. Specifically, we evaluate the proportion of frames containing multiple objects. Videos are filtered out if this proportion exceeds a predetermined threshold. Finally, we discard videos where the motion is too small to be meaningful in the not discarded yet videos, the number of such videos is $N_{stil}$. The overall discard rate $DR$ is defined as
@@ C.1 Discard rate
C.2 Depth Consistency Evaluation
In all the studied experiments, the video camera is orthogonal to the object’s motion and is fixed. This allows us to compute the Physical Invariance and Dynamical scores using only information extracted from 2D pixel space, available for generated videos. The results are presented in Fig. A7, showing that most of the models are reasonably consistent and thus object properties like energy conservation could be also studied using only 2D coordinates.
Figure A7: Average depth consistency for different video generation models across all studied experiments.
C.3 Physically-informed Neural Networks
@@
$L_{\text{total}}=L_{\text{data}}+\lambda L_{\text{physics}},$ (A9)
@@
$L_{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{T}_{i}-T_{i}\|^{2},$ (A10)
@@
$L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{y}}_{j}+g\right\|^{2},$ (A12)
@@
$\ddot{\theta}+\frac{g}{l}\sin\theta=0,$ (A13)
@@
$L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{\theta}}_{j}+\frac{g}{l}\sin(\hat{\theta}_{j})\right\|^{2},$ (A14)
@@
$\text{NMSE}=\frac{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}},$ (A15)
@@
C.4 Physical Invariances
Table A2: Conserved quantities for each physical experiment in an ideal case.
@@ C.4 Physical Invariances
non-holonomic pendulum low resistance energy (approximately)
@@
$\frac{E}{m}=\frac{1}{2}(v_{x}^{2}+v_{y}^{2})+gy=\text{const}$ (A17)
@@
$\frac{p_{x}}{m}=V_{x}=\text{const}$ (A19)
We provide some examples of estimated invariants in Fig. LABEL:fig:falling_ball_analysis
@@
$\ddot{\theta}+\frac{b}{m}\dot{\theta}+\frac{g}{l}\sin\theta=0$ (A20)
@@
$T=2\pi\sqrt{\frac{l}{g}}\sqrt{1-\left(\frac{b}{2m\omega_{0}}\right)^{2}}$ (A22)
@@
Non-holonomic Pendulum
For an ideal non-holonomic pendulum, energy conservation holds since non-holonomic constraints only restrict possible paths without energy dissipation. However, in laboratory conditions, the suspension string’s deformation and friction cause some energy decay. As with the holonomic pendulum case, we measure this energy decay range in real-world videos and use it to evaluate generated videos.
Since both period and pendulum length vary in non-holonomic pendulums, we exclude these metrics from our analysis, keeping only the energy.
C.5 Physical Score Scaling
@@
$C_{\bar{\sigma}}=C_{\sigma}/C_{mean}$ (A23)
@@
$S=\frac{1}{1+\alpha*C_{\bar{\sigma}}}$ (A24)
@@
so it can be neglected that the influence of mean energy/acceleration is near zero. In the experiment, if $|C_{mean}|\geq C_{threshold}$, we calculate the Physical score as defined in Eq. A24. Otherwise, we use the following Eq. A26 that takes the absolute standard deviation rather than the relative standard deviation.
$S=\frac{1}{1+\alpha*C_{\sigma}}$ (A26)
@@
This method has two key limitations. First, the scores are highly sensitive to the choice of time window size. Larger time windows tend to yield lower scores as they encompass more fluctuations in the trajectory. Second, the method may fail to detect unphysical behavior in generated videos where objects remain stationary for long periods. Since we only consider the highest score among all segments, these periods of stillness - which would receive low scores - are effectively ignored in our evaluation. In future work, we plan to introduce a penalty term to specifically address and discourage such unphysical stillness behavior.
Appendix D Video Generative Models Details
@@
At their core, latent video generative models often utilize a combination of a 3D Variational Autoencoder (VAE) [41, 42] to tokenize individual frames, a text encoder, like T5 [43] to encode frames into latent. During training a noisy latent is produced by the forward diffusion process. This latent is then processed by a parametrized model, either a transformer model  [44] or a U-Net [45, 46, 47] resulting in a patchified long sequence of visual tokens, in case of the former type of model.
@@
Depending on the model architecture different input modalities can be handled like text-to-video, image-to-video, text+image-to-video, and sometimes video continuation regimes facilitating both open-domain and controlled generation scenarios [16, 48, 5, 49]. Although some state-of-the-art video generation models adopt an autoregressive framework, predicting frames sequentially based on prior outputs [50, 51, 52], many others utilize non-autoregressive approaches to generate frames simultaneously [16, 48, 53]. In Table A3, we specify the parameters of particular models used in our benchmark, along with it’s architectural design choices. As to faithfully get the best generation outcome we use the best hyperparameters reported for each model. Due to the high API costs-per five seconds video, along with the impressive performance of COSMOS [5] among established benchmarks, we opted to not use any closed-source VGMs like [7, 54]. Still all the open-source alternatives used for our analysis match the performance of most closed-source one across established benchmarks for VGMs.
Model Category Resolution Number of Video Frames Guidance Scale Sampling Steps
CogVideoX single-frame generation 960 x 768 84 6.0 50
Cosmos single-frame generation 1280 × 704 121 7.0 35
LTX-Video single-frame generation 960x736 81 3.0 50
PyramidalFlow single-frame generation 1280 x 768 121 4.0 10
@@
Table A3: Details of video generation models adopted in our benchmark study, including their category, resolution, number of video frames, guidance scale, and sampling steps.
Appendix E Prompts for Video Generation Models
For each experiment, we carefully designed a prompt that describes the physical setup and motion of the experiment being conducted. For example, in the falling ball experiment, the prompt specifies that the ball falls and makes contact with the table below. Similarly, in the projectile experiment, we describe how the ball is launched at a slight upward angle and follows a natural parabolic trajectory. We enhance these prompts using ChatGLM to incorporate more detailed scene descriptions and contextual elements derived from the reference images. All prompts are shown in Table A4,

### v2

@@ Appendix A Dataset
Overall, we conducted a set of 9 core physical experiments, highlighting different physical principles, including:
  1. 1.
    Falling: Objects dropped from rest until they make impact with the surface, used to test uniform gravitational acceleration and energy conservation.
  2. 2.
  3. 3.
    Rolling: A metal can rolling from a slope, with energy conservation.
  4. 4.
    Sliding: A book sliding from a slope, with energy conservation.
  5. 5.
  6. 6.
  7. 7.
    Bouncing: A ball observed from the moment it first impacts the surface until it rebounds and impacts again, testing gravitational acceleration and energy conservation in a more challenging setting.
  8. 8.
    Collision: Two metal balls with the same or different masses collide with each other. One of them is initially stationary and the other one collides with it.
  9. 9.
    Spring: A weight hanging below a vertical spring, perform up and down simple harmonic vibration with energy conservation.
For each system, we recorded multiple times the type of experiment trying to have homogenous videos, while after a few iterations, we varied the initial conditions or configuration parameters. Table 1 below summarizes the number of recordings and configurations for each experiment.
Experiment Videos
Factors of
Variation
Configuration Initial Condition Description
Falling 20 2 Object type and height from which the object was released.
@@
Rolling 15 2 Incline angle of the ramp from which the object was released and object type.
Sliding 10 1 Incline angle of the ramp (slope) from which the object was released.
Collision 12 3 Masses of the colliding objects and their initial velocities before impact.
Spring 7 1 Magnitude of the initial force/impulse used to displace the mass from rest.
Table 1: Summary of the experimental dataset from real-world recorded videos.
For each experiment, we provide representative frames: falling objects in Figure 9; bouncing, projectile motion, holonomic pendulum, and double pendulum in Figure 10; rolling in Figure 11; and sliding, collision, and spring in Figure 12.
Figure 8: Morpheus scores for each experiment on real-world videos.
Falling
For the falling experiment, we started with a standard table tennis orange ball as the simplest object. A mechanical actuator  [^1] Motor Model: T825, Motor serial number: 00362129 was used to hold the ball at a certain height (initial position) and as a release mechanism to control the moment the ball fell free, before making contact with the surface below. Different height levels from the surface were used as initial positions, resulting in trajectories with different lengths (smaller or larger). In addition to the ball, we conducted extra experiments using different everyday objects, namely a plastic whiteboard marker, an adhesive tape roll, and an apple. Unlike actuator controlled experiments, these objects were released directly from the hand of a person at varying initial heights. This setup introduced additional variability in the conditions and the orientation of the object.
@@
Rolling
For this setting, we examined objects that roll down an inclined surface. We used three different objects: a full can, which was sealed, an empty can, and an orange. The slope of the surface was adjustable, allowing us to vary the steepness of the slope. The objects were placed by hand at the top of the slope before being released. Due to different mass distributions and shapes, we had different rolling behaviors across these three objects.
Sliding
For this experiment, we investigated the sliding motion on an inclined surface using a flat book. We varied the slope of the surface between trials. The book was placed by hand at the top of the slope before being released. Depending on the angle of the slope, the trajectories exhibited smooth sliding motion.
Collision
In the collision experiment, we studied collisions between objects of different sizes. Three setups were recoded: A large object collides with a smaller one, two objects of equal size collide with each other, and a small object collides with a larger one. For all cases, the initial velocity of the moving object (leftmost object with the right one at rest) was introduced with a gentle push at random speeds. These settings produced a diverse outcome for the aforementioned experiment, depending on the relative mass of the objects and the initial velocity.
Spring
In this spring experiment, we analyzed the oscillatory motion of a cylindrical metal weight suspended from a vertical spring. The object was initially displaced from its rest position by hand with a small (but random) force and then released. The amplitude of oscillation was defined by this displacement and the restoring force led the object to move in a vertical periodic movement until equilibrium. Initial force was the main reason for the motion, with no actuator being employed, and gradually over time decayed due to damping effects.
Figure 9: Representative frames from the falling experiments in the Morpheus benchmark: ball, apple, tape, and marker.
Figure 10: Representative frames from four experiments in the Morpheus benchmark: bouncing ball, projectile motion, holonomic pendulum, and double pendulum.
Figure 11: Representative frames of the rolling experiments in the Morpheus benchmark: an empty can, a full can, and an orange, each released by hand on an inclined surface of varying slope.
Figure 12: Representative frames from three experiments in the Morpheus benchmark: sliding, collision (equal-sized objects), and spring.
Appendix B Augmentation with Cosmos-Transfer1
We augment a subset of experiments with style transfer to diversify object appearances while preserving underlying physical dynamics, creating initializations more representative of typical VGM training data. Specifically, we apply COSMOS-Transfer to five experiment types—Falling ball, Projectile, Holonomic pendulum, Rolling, and Sliding. For each experiment type, we select representative videos and generate style-transferred variants that alter object semantics and appearance while retaining motion consistency. This yields three transferred variants per original video (two for Holonomic pendulum), providing diverse yet physically plausible conditioning scenarios. Per experiment type, this produces $3\times 3=9$ augmented conditioning scenarios for most experiments, and $3\times 2=6$ for Holonomic pendulum.
The transfer process uses per-frame object masks from SAM2 Ravi et al. (2024), reusing the same masks from our trajectory-extraction pipeline to isolate only the object under study. We provide concise, object-focused prompts describing the desired replacement. All transfers undergo manual screening for visual and physical plausibility. When artifacts such as camera motion, temporal drift, hallucinations, shape misalignment, or mask leakage are observed, we enable Canny-edge guidance in COSMOS-Transfer with an edge-conditioning weight of 0.5, with the object mask now accounting for the other 0.5 conditioning weight. This additional constraint preserves scene structure while still allowing complete style transformation.
Figure 13 shows representative examples comparing the first frames of original videos with their style-transferred variants. These augmentations effectively multiply our conditioning scenarios: each original video contributes multiple physically consistent variants, expanding the diversity of initial conditions for VGM evaluation.
Figure 13: COSMOS-Transfer style augmentations for physics experiments. Each row displays the original experiment (left) with its style-transferred variants (right), showing how object appearance can be altered while preserving physical dynamics. The transfers provide diverse yet physically consistent initializations for VGM evaluation across five fundamental physics scenarios.
Appendix C Velocity and acceleration estimation
@@
$\begin{bmatrix}v\\ b\end{bmatrix}=\left(A^{T}A\right)^{-1}A^{T}x$ (3)
@@
$A=\begin{bmatrix}t_{1}&1\\ t_{2}&1\\ \vdots&\vdots\\ t_{w}&1\end{bmatrix}$ (4)
@@
$v_{\text{final}}=\alpha v_{\text{regression}}+(1-\alpha)v_{\text{central}}$ (5)
Here $\alpha=0.7$, indicating greater confidence in the regression method. Finally, we apply Savitzky-Golay filtering for smoothing (Luo et al., 2005). This step effectively removes high-frequency noise from velocity calculations.
$v_{\text{smoothed}}=\text{SG}(v_{\text{final}},\text{window},3)$ (6)
@@
$v(t)=\text{SG}\left(\alpha v_{\text{regression}}(t)+(1-\alpha)v_{\text{central}}(t),w,3\right)$ (7)
where w is the window size (odd number for symmetry); $\alpha=0.7$ is the weighting coefficient; SG represents Savitzky-Golay filter of order 3; Regression window range: [tw/2,t+w/2][t-w/2,t+w/2].
For the acceleration, we first calculate the acceleration using the central difference. For $1\leq i\leq N-2$:
$a_{i}=\frac{v_{i+1}-v_{i-1}}{t_{i+1}-t_{i-1}}$ (8)
@@
Appendix D Evaluation metrics
Benchmark Real-world ground-truth Quantitative evaluation Initial condition grounding Physical laws evaluation
VideoCon-Physics  (only text)
VAMP  (only text)
PhyGenBench  (only text)
Kang et al. (2024)  (1 or 3 frames)
COSMOS  (image + video)
Physics-IQ  (image + video)
Morpheus (ours)  (image + video)
Table 2: Comparison of physics-based video understanding benchmarks. Our benchmark is the first to use real physical laws for evaluation. Symbols: = supported, = not supported
D.1 Discard rate
We generate $N_{total}$ videos for each type of experiment. Among these videos, we discard those that do not meet our quality standards, following a three-stage filtering out. First, we discard videos where object are disappearing from the videos the number of such videos is $N_{disappear}$. Second, we analyze the number of objects in each video and discard videos that do not maintain a consistent object count in the not discarded yet videos. For this purpose we employ DEVA tracking (Cheng et al., 2023) built on top of Grounded SAM (Ren et al., 2024) (with object names from the prompt as Grounding DINO (Liu et al., 2025) query) for consistent open-vocabulary prediction of 2D object masks. We denote the number of discarded videos in this step as $N_{duplicate}$. Specifically, we evaluate the proportion of frames containing multiple objects. Videos are filtered out if this proportion exceeds a predetermined threshold. Finally, we discard videos where the motion is too small to be meaningful in the not discarded yet videos, the number of such videos is $N_{stil}$. The overall discard rate $DR$ is defined as
@@ D.1 Discard rate
D.2 Depth Consistency Evaluation
In all the studied experiments, the video camera is orthogonal to the object’s motion and is fixed. This allows us to compute the Physical Invariance and Dynamical scores using only information extracted from 2D pixel space, available for generated videos. The results are presented in Fig. 14, showing that most of the models are reasonably consistent and thus object properties like energy conservation could be also studied using only 2D coordinates. Proprietary Veo3 and Veo3-fast (Veo-Team et al., 2024), which were also initially investigated, were excluded from the final analysis due to budget limits to compute a similar batch as other models, however, we plan to regularly update our leaderboard with state-of-the-art models’s evaluation.
D.3 Physically-informed Neural Networks
@@
$L_{\text{total}}=L_{\text{data}}+\lambda L_{\text{physics}},$ (9)
@@
$L_{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{T}_{i}-T_{i}\|^{2},$ (10)
@@
$L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{y}}_{j}+g\right\|^{2},$ (12)
@@
$\ddot{\theta}+\frac{g}{l}\sin\theta=0,$ (13)
@@
Figure 14: Average depth consistency for different video generation models across all studied experiments.
@@
$L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{\theta}}_{j}+\frac{g}{l}\sin(\hat{\theta}_{j})\right\|^{2},$ (14)
@@
$\text{NMSE}=\frac{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}},$ (15)
@@
D.4 Physical Invariances
Table 3: Conserved quantities for each physical experiment in an ideal case.
@@ D.4 Physical Invariances
sliding uniform surface forces acceleration
rolling uniform surface forces acceleration
elastic collision perfect elasticity total energy, linear momentum
spring-mass system ideal Hookean spring period
@@
$\frac{E}{m}=\frac{1}{2}(v_{x}^{2}+v_{y}^{2})+gy=\text{const}$ (17)
@@
$\frac{p_{x}}{m}=V_{x}=\text{const}$ (19)
We provide some examples of estimated invariants in Fig. 26(a).
@@
$\ddot{\theta}+\frac{b}{m}\dot{\theta}+\frac{g}{l}\sin\theta=0$ (20)
@@
$T=2\pi\sqrt{\frac{l}{g}}\sqrt{1-\left(\frac{b}{2m\omega_{0}}\right)^{2}}$ (22)
@@
D.5 Physical Score Scaling
@@
$C_{\bar{\sigma}}=C_{\sigma}/C_{mean}$ (23)
@@
$S=\frac{1}{1+\alpha*C_{\bar{\sigma}}}$ (24)
@@
so it can be neglected that the influence of mean energy/acceleration is near zero. In the experiment, if $|C_{mean}|\geq C_{threshold}$, we calculate the Physical score as defined in Eq. 24. Otherwise, we use the following Eq. 26 that takes the absolute standard deviation rather than the relative standard deviation.
$S=\frac{1}{1+\alpha*C_{\sigma}}$ (26)
@@
This method has two key limitations. First, the scores are highly sensitive to the choice of time window size. Larger time windows tend to yield lower scores as they encompass more fluctuations in the trajectory. Thus, we kept time window constant between evaluating different model generations. Second, the method may fail to detect unphysical behavior in generated videos where objects remain stationary for long periods; however, this is addressed by discarding videos due to stillness.
Appendix E Extended Limitations
The main focus of our benchmark is towards a set of Newtonian scenarios, which take place under a controlled environment and static camera conditions. While this enables clean invariants and can be reproducible, it limits a full scale physical coverage. Additionally, ambiguity is introduced (e.g. distortion of lens, unknown scale, etc.) because all of our estimates come from 2D pixel trajectories without any camera calibration. At the same time, we rely on assumptions such as negligible air resistance and friction. Unmodeled forces, such as air drag, friction and rotational kinetic energy could cause violations, which do not reflect generations that are wrong but rather evaluator mismatch. We perform trajectory extraction from generated videos using fixed first/last frame or short multi-frame conditioning that cannot specify initial conditions fully, such as the mass of an object or the coefficient of the spring. Moreover, a physically plausible generation can be penalized heavily when some of the assumptions are violated, leading to misleading results. Similarly, one single invariant can look really "good", but the generated video could violate multiple laws, making it physically implausible.
Errors in segmentation and tracking introduce noise. Additionally, we are only considering short clips, and the model might drift away from physical laws over a longer temporal horizon, which our current metrics would not be capable of capturing. Finally, it should be noted that while our scores remain proxies rather than direct measurements of underlying physical plausibility, that would require falsifiable and controllable VGM generation. Prior benchmarks that rely on judgments from humans or VLMs are prone to hallucinations and often miss slight inconsistencies. In contrast, our scores avoid these issues but still need to be interpreted jointly to give a much more reliable picture. With that in mind, Morpheus should be treated as a focused and reproducible stress test for core physical laws and not a benchmark that can test and evaluate all physical dynamics that might take place in a single video.
Appendix F Video Generative Models Details
@@
At their core, latent video generative models often utilize a combination of a 3D Variational Autoencoder (VAE) (Kingma, 2013; Lin et al., 2024) to tokenize individual frames, a text encoder, like T5 (Raffel et al., 2020) to encode frames into latent. During training a noisy latent is produced by the forward diffusion process. This latent is then processed by a parametrized model, either a transformer model  (Yang et al., 2023) or a U-Net (Ronneberger et al., 2015; Zhou et al., 2022; Bar-Tal et al., 2024) resulting in a patchified long sequence of visual tokens, in case of the former type of model.
@@
Depending on the model architecture, different input modalities can be handled like text-to-video, image-to-video, text+image-to-video, and sometimes video continuation regimes facilitating both open-domain and controlled generation scenarios (Yang et al., 2024b; Kong et al., 2024; Agarwal et al., 2025; Brooks et al., 2024b). Although some state-of-the-art video generation models adopt an autoregressive framework, predicting frames sequentially based on prior outputs (Weng et al., 2024; Deng et al., 2024; Weissenborn et al., 2020), many others utilize non-autoregressive approaches to generate frames simultaneously (Yang et al., 2024b; Kong et al., 2024; Xing et al., 2025). In Table 4, we specify the parameters of particular models used in our benchmark, along with it’s architectural design choices. As to faithfully get the best generation outcome we use the best hyperparameters reported for each model. Due to the high API costs-per five seconds video, along with the impressive performance of COSMOS (Agarwal et al., 2025) among established benchmarks, we opted to not use any closed-source VGMs like (Brooks et al., 2024a; Kling AI, 2024). Still all the open-source alternatives used for our analysis match the performance of most closed-source one across established benchmarks for VGMs.
Model [Params.] Resolution
Number of
Video Frames
Guidance
Scale
Sampling
Steps
CogVideoX [5B] 960 x 768 84 6.0 50
Cosmos-Predict 1 [14B] 1280 × 704 121 7.0 35
Cosmos-Predict 2 [14B] 1280 × 704 93 7.0 35
WAN2.1 [14B] 1280 × 720 121 5.0 40
LTX-Video [13B] 960x736 81 3.0 50
PyramidalFlow[12B] 1280 x 768 121 4.0 10
@@
Table 4: Details of video generation models adopted in our benchmark study, including their resolution, number of video frames, guidance scale, and sampling steps.
@@
Projectile A single, small 3D-printed ball, dark gray in color, is launched from a plastic, small-scale ramp with a slight upward angle. The ball follows a natural, smooth, arcing trajectory upward and then downward, continuing along that arc until it exits the right side of the video frame. The video should accurately simulate the ball’s motion under standard earth gravity, showing a clear parabolic arc. The video should emphasize a smooth and realistic physics-based movement of the ball without any sudden changes in speed. The ball should not bounce or collide with any objects in the scene. Fixed camera view, no camera movement. In a meticulously crafted scene, a solitary, dark gray 3D-printed ball, with its sleek, spherical form, is propelled from a plastic ramp that slopes gently upward. The ball, weighing a mere fraction of a kilogram, is captured in high-definition, showcasing every nuance of its motion. As it leaves the ramp’s edge, the ball arcs gracefully into the air, its trajectory a perfect parabola that mirrors the laws of physics under standard earth gravity. The video’s frame follows the ball’s smooth ascent and descent, highlighting the ball’s consistent speed and the absence of any sudden accelerations or decelerations. The scene remains unobstructed, ensuring that the ball’s journey is uninterrupted by any external forces, save for the pull of gravity, resulting in a visually stunning and scientifically accurate demonstration of a parabolic motion. Fixed camera view, no camera movement.
Rolling Can An empty soda can rolls down a wooden surface. Experiment carried out in a laboratory controlled environment. Fixed camera view, no camera movement. An empty aluminum red soda can rolls steadily down an inclined wooden board in a controlled laboratory environment. This motion should be depicted with physical realism, accurately simulating the key dynamics: the can accelerates under gravity, with its combined rotational and translational movement appearing authentic and governed by the frictional interaction with the wooden surface, ensuring a physically plausible rolling movement. Fixed camera view, no camera movement.
Sliding Book A book slides down a wooden surface. Experiment carried out in a laboratory controlled environment. Fixed camera view, no camera movement. A book slides steadily down an inclined wooden board in a controlled laboratory environment. This motion should be depicted with physical realism, accurately simulating the key dynamics: the book accelerates due to the component of gravity acting along the incline, opposed by kinetic friction from the wooden surface. Its movement should be purely translational, maintaining consistent contact with the board and without any significant rotation or tumbling, ensuring a physically plausible descent. Fixed camera view, no camera movement.
Holonomic Pendulum A single pendulum moving retrogressively back and forth. At the bottom of a pendulum, there is a ball attached to it. The pendulum is holonomic. Fixed camera view, no camera movement. A pendulum with a spherical ball attached swings back and forth in a controlled manner, its motion captured in a moment of retrograde swing. The pendulum’s arm, likely made of metal, extends horizontally from a stand, connected to a pivot point that allows for rotational movement. The ball, positioned at the lower end of the pendulum, appears to be in motion, indicating the pendulum’s swing. The environment suggests a laboratory or testing setting, with a backdrop of technical apparatus and equipment, and the lighting is artificial, casting a uniform glow over the scene. The pendulum’s movement, while currently in a retrogressive swing, could potentially change direction, continuing its oscillatory motion. Fixed camera view, no camera movement.
Table 5: Base and enhanced textual prompts used for video generation experiments: Falling Ball, Projectile, Rolling Can, Sliding Book, Holonomic Pendulum. Enhanced prompts are generated using corresponding upsampler VLMs such as ChatGLM (Zeng et al., 2024) and incorporate more detailed scene descriptions and contextual elements derived from the reference images. Slight modifications to these prompts were made in case of different variants such as object type.
Experiment Name Base Prompt Enhanced Prompt